Is saving necessarily invested or not.
How did Keynes resolve this contradiction
1
Expert's answer
2020-05-20T09:40:24-0400

Keynes had come up with the theory that savings and investment are equal. However, that is not always the case. This contradiction was resolved by Keynes through; the assumption that a country produces two kinds of goods- investment (I) and consumer (C) goods. Income gets spent though buying of both goods, hence, national income is sum of investment and consumption demand implying that some part of income is consumption while the rest is saved; C + I = Y = C + S hence I.=S. The calculation proposes that savings and investment are equal at equilibrium showing the contradiction. To resolve the contradiction, actual saving, SA, actual investment, IA, planes investment IP and planned savings, SP are introduced. Thus, SA = IA and SP = IP.  but only at equilibrium is are they equal and unequal at disequilibrium income level. Therefore, investment and saving are equal through the ex-post sense interpretation. Planned investment and saving behave the same as actual saving and investment at equilibrium and disequilibrium. Keynes asserts that investors and savers are two distinct individuals in the community and behave differently. This means that motives of both individuals are different. Hence, there is insufficient reason to believe that what a household desires to save is the exact amount an investor may desire to invest. There is inadequate reason why desires savings should equate desired investment at a random choice of income, however, when they are unequal in the two-sector economy, national income will shift until they are both equal.